Webbie10000:

A skyscraper in London, England, has a roof in the shape of a square pyramid. Its approximate dimensions are shown in the net. What is the surface area of the skyscraper's roof, including the base? Enter your answer in the box. m²

TheSmartOne:

Where are the dimensions? Post a screenshot.

Webbie10000:

1 attachment
Pixel:

OOf i remember surface area

Webbie10000:

someone pls help i am struggling

TheSmartOne:

What's the area of one of the triangles. Area for a triangle = 1/2 *base * height

Webbie10000:



AnimeGhoul8863:

For this question, find the area of each individual figure and then add all of them together. Let's start with the triangles. To find the area of a triangle, use the formula A = bh (b = base and h = height). A = bh A = (30)(40) A = (1200) A =  A = 600m² There are 4 triangles, so multiply the area of one triangle by 4 to find the area of all 4 triangles. (600)(4) = 2400m² Now we can find the area of the square. To find the area of a square, simply multiply the length x width (or square one side). A = l x w A = (30)(30) A = 900m² Add together the area of the triangles and the area of the square to get the total area: 2400 + 900 =  3300m²

Webbie10000:

thanks so the answer is 3300m2

BenLindquist:

technically no direct answers but who cares

AnimeGhoul8863:

its not direct its showing her how to do it

BenLindquist:

still gave the answer. techinally

Webbie10000:

yeah thanks for showing me how to do it

TheSmartOne:

You didn't guide them to the answer. All they learned from this was the answer

BenLindquist:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TheSmartOne You didn't guide them to the answer. All they learned from this was the answer \(\color{#0cbb34}{\text{End of Quote}}\) thank you

AnimeGhoul8863:

no they learn the answer and how to do it bc they see the steps

BenLindquist:

the questioner is supposed to tell you how they got to the answer gg

Webbie10000:

A rectangular prism has a length of 7 cm, a width of 4 cm, and a height of 2 and 1/2 cm. The prism is filled with cubes that have edge lengths of 1/2 cm. How many cubes are needed to fill the rectangular prism? Enter your answer in the box. To fill the rectangular prism, cubes are needed.

Webbie10000:

help

Ultrilliam:

Please post it as a new question, thank you.

TheSmartOne:

Make a new post for a new question

BenLindquist:

^ and ^^

AnimeGhoul8863:

BEN GET MY FACE OFF NOW!!!!!!!!!!!!!!!!!!!!

Ultrilliam:

That might hurt anime... are you sure?

TheSmartOne:

😂😂

AnimeGhoul8863:

-_- ultri im not in the mood and u know why NOW MAKE HIM GET MY FACE OFF HIS PFP

BenLindquist:

😂😂

Webbie10000:

i sent it in another question

BenLindquist:

nice on jay